Brief Exercise 9-6

Rosco Taxi Service uses the units-of-activity method in computing depreciation on its taxicabs. Each cab is expected to be driven 151,923 miles. Taxi no. 10 cost $40,000 and is expected to have a salvage value of $500. Taxi no. 10 is driven 34,480 miles in year 1 and 20,670 miles in year 2.

Calculate depreciation expense per mile using unit of activity method. (Round answer to 2 decimal places, e.g. 0.50.)

Compute the depreciation for each year. (Round answers to 2 decimal places, e.g. 2,125.50.)
Answer #1

Depreciation expense per mile

= (Cost - Salvage value)/Total miles to be driven

= (40,000-500)/151,923

= 0.26 per mile

Depreciation for

Year 1 = (34,480*0.26) = 8964.80

Year 2 = (20,670*0.26) = 5374.20